Aug/Sept 2013 LUCIFER Recalibration after Installation on the Telescope

Critical Tasks before Science can start
  • Basic MOS and instrument functional checkout
  • Internal focus check
  • Field centering and pupil alignment
  • Check of Flexure Compensation
  • Check of Grating Tilt calibration
  • Orientation on sky (LeftZP)
  • Find Center of Rotation
  • (Recalibrate AGw transform)
  • (Verify new AGw transform)
  • Run full spectroscopic observation on std star

Executive Summary Block
DAYTIME (Closed Dome) Tests . . . . . . . . . . . . .
Test Status Test Status
LUCI1 re-installation on telescope. done FITS header check pending
MOS tests done Field centering and Pupil alignment pending
General function tests done FlexComp - Calibrate pending
Internal Focus determination: done Grating Tilt (Lambda-V) Calibration pending
General reproducibility tests pending FlexComp - Verify pending
Spectral Alignment Check: pending LUCI Aux cryostat - Xchng 04 Feb pending

NIGHTTIME (On Sky) Tests . . . . . . . . . . .
Test Status Test Status
Spectral Alignment Check: pending LUCI Aux cryostat - Xchng 04 Feb pending
Twilight Flats (IFF photometric) pending PM: Field illumination pending
Orientation on sky pending PM: Specphot tests pending
Image scale, all cameras: pending Ghosts, scattered light pending
Find Center of Rotation pending Run a complete science observation pending
Calibration of AGw transform pending LUCI-to-AGw Focus Measurement pending
Field Aberration Measurement pending    
Verification of AGw transform pending    
PM: Photometric ZPs pending    

DAYTIME (Closed Dome) Tests . . . . . . . . . . . . .
  • LUCI1 re-installation on telescope.
  • MOS tests
  • General function tests
  • Internal Focus determination:
  • General reproducibility tests
  • Spectral Alignment Check:
  • FITS header check
  • Field centering and Pupil alignment
  • FlexComp - Calibrate
  • Grating Tilt (Lambda-V) Calibration
  • FlexComp - Verify
  • LUCI Aux cryostat - Mask exchange scheduled for 04 Feb

NIGHTTIME (On Sky) Tests . . . . . . . . . . .
  • Twilight Flats (IFF photometric)
  • Orientation on sky
  • Image scale, all cameras:
  • Find Center of Rotation
  • Calibration of AGw transform
  • Field Aberration Measurement
  • Verification of AGw transform
  • PM: Photometric ZPs
  • PM: Field illumination
  • PM: Specphot tests
  • Ghosts, scattered light
  • Run a complete science observation
  • LUCI-to-AGw Focus Measurement
  • Collimation model check
  • Pointing model check


On-telescope day-time (closed dome) calibrations:

LUCI1 re-installation on telescope. Duration 1.0d
Once done, the following are prerequisites before continuing with the recalibrations
  • Verify MOS GUI to make sure masks are locked!
  • Wait for LUCI to return to thermal equilibrium - Detector stable @77K (verify)
  • Switch on ROE, wait until stable (= time ROE was off, max 2hr)
  • Temps and pressures all look nominal

Backup all configuration files

MOS tests Duration: 4h
Responsible: PB
Telescope: @Zenith, Rotator enabled, Lights can be ON

> Basic MOS checkout:
Tests not detailed. Peter's steps to complete the MOS checkout should be documented for future reference.

done by PBuschkamp

General function tests Duration: 4h
Comments: Use ONLY engineering panels, Main Instrument GUI only usable AFTER Peter has checked the MOS
Telescope: @Zenith, Rotator enabled, Lights can be ON

> Basic instrument function tests:
  • Stop SW / Re-start SW (optional, restarted in move to telescope)
  • Repeat at ROT ~ 341, 221, 101 degrees the following:
    • Camera Wheel: Init; move several times between N1.80 & N3.75 camera
    • FW 1 & 2: init, move both wheels front/backwards movements
    • Grating Unit: init, select all gratings/mirror
    • Grating Tilt: set voltages 0V, +-1V
    • Detector focus: move steps to reach neglim & poslim, note # of steps
    • PV: init and move in/out

done 
20130901
0001 - 0002 1x 4.0s Dark O2DCR  -> RN = 12.75e-, looks oddly noisy
0003 - 0004 1x12.0s Dark MER10  -> RN =  8.35e-, --"--
0052 - 0053 1x 4.0s DCR Dark  -> RN = 12.57e-, looks better
0054 - 0055 1x12.0s MER Dark  -> RN =  7.75e-, looks much better
20130902
0001 - 0002 1x 4.0s DCR Dark  RN = 12.53e- 
0003 - 0004 1x12.0s MER Dark  RN = 7.73e- 

Internal Focus determination: Duration: 0h
Telescope: @Zenith, Rotator @ 341, Lights can be ON
LUCI: Optic Sieve mask

> Detector stage focus calibration:
  • See ~engineer/TemplateScripts/focusN180.txt and focusN375.txt
  • adjust step range if needed
  • Insert best focus values for both cameras into detector unit GUI config
  • Check tilt of detector w.r.t. focal plane in these data
    • Uniformity of out-of-focus images at +extreme of focus curve data

done

0005        1x 4.0s dark for focus
0006 - 0014 DetFocus_N375.img  best focus @3.81 --> 358 steps (2nd coarse ring around center)
                               best focus @4.83 --> 368 steps (4th coarse ring around center)

0015 - 0021 DetFocus_N180.img  best focus @3.50 --> 368 steps (13x9 coarse rectangle)
updates made to config files

To update:
  • Open FOCUS engineering GUI, Edit configuration
  • N375
    • N_3_75_REFERENCE_OFFSET_FROM_NEGLIM change from 410 to 370
    • N_3_75_REFERENCE_OFFSET_FROM_POSLIM change from -220 to -260
    • Yes, these values are the same as for the N180 data!
  • N180
    • No change from current configuration
  • Click "OK" to save changes
  • Initialize and check that the new values are used.

General reproducibility tests Duration: ?h
Telescope: @Zenith, Rotator enabled, Lights OFF in dome
LUCI: Calib unit in beam, Spectro Sieve mask.
Comments:
  • move ~5x between the corresponding units
  • Repeat at 3-4 rotator angles (90deg separation)
  • Take image after each move
  • Measure reproducibility

Reproducibility tests:
  • Grating Unit
    • Use arclamp (or arclamp + flatlamp to see continuum?)
    • only moves grating unit between measurements
    • All gratings + N3.75 combos, see table below
    • Closed tilt loop on grating (request nominal CWL)
    • Move between mirror <-> grating 5x
    • take image at each position
  • Cameras
    • Use flatfield lamp or thermal background
    • Only moves camera wheel between measurements
    • Grating Unit fixed on mirror position
    • Move between N1.80 <-> N3.75 5x
    • take image at each position
  • Calculate rms positioning for each set of data

done 

N3.75
ROT    rmsX    rmsY  
341    0.096   0.122
221    
101    

N1.80
ROT    rmsX    rmsY  
341    0.036   0.141
221    
101    

General function tests Duration: 4h
Comments: Use ONLY engineering panels, Main Instrument GUI only usable AFTER Peter has checked the MOS
Telescope: @Zenith, Rotator enabled, Lights can be ON

> Basic instrument function tests:
  • Stop SW / Re-start SW (optional, restarted in move to telescope)
  • Repeat at ROT ~ 341, 221, 101 degrees the following:
    • Camera Wheel: Init; move several times between N1.80 & N3.75 camera
    • FW 1 & 2: init, move both wheels front/backwards movements
    • Grating Unit: init, select all gratings/mirror
    • Grating Tilt: set voltages 0V, +-1V
    • Detector focus: move steps to reach neglim & poslim, note # of steps
    • PV: init and move in/out

Internal Focus determination: Duration: 0h
Telescope: @Zenith, Rotator @ 341, Lights can be ON
LUCI: Optic Sieve mask

> Detector stage focus calibration:
  • See ~engineer/TemplateScripts/focusN180.txt and focusN375.txt
  • adjust step range if needed
  • Insert best focus values for both cameras into detector unit GUI config
  • Check tilt of detector w.r.t. focal plane in these data
    • Uniformity of out-of-focus images at +extreme of focus curve data

To update:
  • Open FOCUS engineering GUI, Edit configuration
  • N375
    • N_3_75_REFERENCE_OFFSET_FROM_NEGLIM change from 410 to 370
    • N_3_75_REFERENCE_OFFSET_FROM_POSLIM change from -220 to -260
    • Yes, these values are the same as for the N180 data!
  • N180
    • No change from current configuration
  • Click "OK" to save changes
  • Initialize and check that the new values are used.

General reproducibility tests Duration: ?h
Telescope: @Zenith, Rotator enabled, Lights OFF in dome
LUCI: Calib unit in beam, Spectro Sieve mask.
Comments:
  • move ~5x between the corresponding units
  • Repeat at 3-4 rotator angles (90deg separation)
  • Take image after each move
  • Measure reproducibility

Reproducibility tests:
  • Grating Unit
    • Use arclamp (or arclamp + flatlamp to see continuum?)
    • only moves grating unit between measurements
    • All gratings + N3.75 combos, see table below
    • Closed tilt loop on grating (request nominal CWL)
    • Move between mirror <-> grating 5x
    • take image at each position
  • Cameras
    • Use flatfield lamp or thermal background
    • Only moves camera wheel between measurements
    • Grating Unit fixed on mirror position
    • Move between N1.80 <-> N3.75 5x
    • take image at each position
  • Calculate rms positioning for each set of data

Spectral Alignment Check: Duration: 0h
Use flatfield lamp for these tests.
  • For each grating+camera combo in table below:
    • Make sure grating tilt stabilization is ON
    • Set to grating to nominal CWL, take dispersed image
    • Set to mirror, take image
    • Only 1 pair of exposures needed for each combination
  • Measure spectrum tilt along the rows
  • Measure vertical (y) displacement between image and spectrum

Define:
  • Ytilt = Ymax - Ymin along dispersed spectrum near Y~1024
  • Yoffset = Yimg - Yspec for a centered pinhole @ X~1024

FITS header check Duration: 0h (offline)
Check the FITS headers for any 'strangeness' and note any other issues

From April 2012 tests: Need to update (for N3.75 camera):
  • PIXSCALE = 0.11782
  • CRPIX1 = 1010.58
  • CRPIX1 = 1047.14
  • SECPIX1 = -0.11782
  • SECPIX2 = 0.11782
  • CDELT1 = 3.272778e-05
  • CDELT2 = 3.272778e-05
  • CD matrix as appropriate
Note apparently no status variable for flexure compensation (ON/OFF)!
  • M4M*POS not an indicator (FC ON but these numbers are not changing)
  • Mirror 4 (M4) not the unit that does flexure compensation?

Field centering and Pupil alignment Duration: 5h
Telescope: @Zenith, Rotator enabled
LUCI: Imaging, N3.75, K band

> Field centering procedure using FM4
  • Take an image and save it
  • Measure gaps at edges of image
  • FM4: move the mirror in steps to center, notes:
    • Enlarge CompensationMirrorGUI to full screen, use right side of it.
    • M1 -> y, M2-> x; + x steps -> moves image to the right, +y to the top
    • 100 steps corresponds to approximately to 1pix
  • Take images with both cameras, verify centering

> Pupil alignment procedure:
  • Take image to show present alignment of the pupil
  • Adjust FM1 + FM4 (WS: document procedure for future reference)?
  • Repeat until alignment is acceptable

Flexure Compensation - calibration data Duration: 8h
Telescope: EL and Rotator enabled, Lights OFF in dome
LUCI: Calib unit in beam
Scripts located in ~observer/LBTO/FlexComp_Calib/

> Flexure compensation calibration:
  • Exposure times need to be checked (4s for img, 20s for spec)!
  • MANUALLY TURN FLEXURE COMPENSATION OFF to measure!
  • Best sequence:
    • one config
    • all ROT positions (script does this, ROT starts/ends @ -88d)
    • 3 EL's: 90, 55, 20 degrees (order does not matter)
    • next config
  • Use AR lamp for the grating measurements!
  • Process data following ~/LBTO/Flexures/FlexureCheck_Luci1.pdf
  • Select same source in all three datasets for one config!
  • Rename object_0001.tbl like: N180_G210_E90.tbl (see rename.sh)
  • Send renamed tables to Walter

Camera Filter Grating Mask Lamp Script Name Status
N3.75 clear K Mirror Spectro Sieve Thermal FC_N375_mirror.cal  
N1.80 clear K Mirror Spectro Sieve Thermal FC_N180_mirror.cal  

Turn on Ar lamp manually!

Camera Filter Grating Mask Lamp Script Name Status
N1.80 HKspec 200_H+K Spectro Sieve Ar FC_N180_200HK.cal  
N1.80 clear J 210_zJHK Spectro Sieve Ar FC_N180_210zJHK.cal  
N1.80 clear Ks 150_Ks Spectro Sieve Ar FC_N180_150Ks.cal  
             
N3.75 clear Ks 150_Ks Spectro Sieve Ar FC_N375_150Ks.cal  
N3.75 clear J 210_zJHK Spectro Sieve Ar FC_N375_210zJHK.cal  
N3.75 HKspec 200_H+K Spectro Sieve Ar FC_N375_200HK.cal  

Grating Tilt (Lambda-V) Calibration Duration: 2d
Telescope: @Zenith, Rotator enabled, Lights OFF in dome
LUCI: Calib unit in beam, LS0.75_450um slit, grating tilt loop closed.

Spectra: grat*
Throughslit-images: ls*
Done for N1.80-camera only

Test table:
Camera Filter Grating Mask Lamp Voltages to test
N1.80 HKspec 200_H+K LS0.75_450um Ar/Xe/Ne 3 different
N1.80 clear z 210_zJHK LS0.75_450um Ar/Xe/Ne 3 different
N1.80 clear J 210_zJHK LS0.75_450um Ar/Xe/Ne 3 different
N1.80 clear H 210_zJHK LS0.75_450um Ar/Xe/Ne 3 different
N1.80 clear K 210_zJHK LS0.75_450um Ar/Xe/Ne 3 different
N1.80 clear Ks 150_Ks LS0.75_450um Ar/Xe/Ne 3 different
N1.80 zJspec 200_H+K LS0.75_450um Ar  

First results (measured position of line at the detector in x, input was lambda of line):
If all is OK, lines should always be at the same position within a few pixels.

Flexure Compensation - verification data Duration: 4h
Telescope: EL and Rotator enabled, Lights OFF in dome
LUCI: Calib unit in beam (OK for imaging too, same temp as dome)
Scripts located in /home/observer/LBTO/FlexComp_Verify/
Comment: Verification, only a subset needed
Prereq: Flexure calibrations in place, flexure compensation ON!

> Flexure compensation verification:
  • MANUALLY TURN FLEXURE COMPENSATION ON to verify FC!
  • MCS GUI -> ROTATORS -> MOTION CONTROL, set max speed to 1 deg/sec
  • Set EL, send LUCI rotator to -69 degrees to start
  • Best sequence:
    • one config
    • all ROT positions (script does this, ROT starts/ends @ -69d)
    • 2 EL's: 38 and 72 degrees (mid-points of measurements)
    • next config
  • Process data following ~/LBTO/Flexures/FlexureCheck_Luci1.pdf
  • Copy results files into ~/LBTO/Flexures/FlexComp_Verify
  • MCS GUI -> ROTATORS -> MOTION CONTROL, set max speed to 5 deg/sec
  • Send resulting data to Walter

Camera Filter Grating Mask Lamp Script Name
N3.75 clear K Mirror Spectro Sieve Thermal FC_N375_mirror.cal

Turn on Ar lamp manually!

Camera Filter Grating Mask Lamp Script Name
N1.80 HKspec 200_H+K Spectro Sieve Ar FC_N180_200HK.cal
N1.80 clear J 210_zJHK Spectro Sieve Ar FC_N180_210zJHK.cal

Other notes: subtract a dark or a lamp-off before processing. Should see <~2 pix residuals.

On-telescope night-time on-sky calibrations:

Note:
  • PM = Photometric

Twilight Flats : Duration: 0h (Twilight)
Evenings/Mornings when photometric:
Pre-requisite: clear skies
Twilight Flatfields

  • Recommend keeping the twilight flatfield counts below ~15000 ADU if at all possible
  • Broadband flats need to start immediately after sunset
  • Adjust focus through BrG before starting
  • Random dither 10x if possible

Orientation on sky: Duration: 2h
Pre-requisite: Orientation done before AGw transform calibration!
Pre-requisite: field with good astrometric coords, good seeing (<1")
Note: Can be done on Persson stds during PM night
Note: Will update further after some MOS acquisitions

> Orientation data
  • Configure for imaging with N3.75, Ks
  • Preset to field (Persson standards, any low-lat field)
  • Take image pair, script: /home/observer/LBTO/SciencePrep/imagepair
  • Repeat for ~5 fields
  • Derive PA offset, update LEFTZEROPOINT
  • Update PCSInstrument.conf and RESTART PCS!
  • Repeat observations for ~5 fields
  • Check that orientation is OK

Note: systematic residuals in the rotator calibration will show as a consistent rotator offset in MOS mask acquisitions that can be "mined" from the event log after science blocks and the rotator ZPs updated. The systematic dX and dY offsets from MOS alignments should also be a measure of the difference between the pointing origin and the mechanical mask center. This should be true for MODS as well as LUCI.

Image scale, all cameras: Duration: 0h
Note: Can use the orientation data (scale is a byproduct)
N3.75 scale expected to change because of lens swap
Pre-requisite: fields with good astrometric coords, good seeing (<1")
Should approximately be N3.75=0.118"/pix, N1.80=0.249"/pix

Find Center of Rotation Duration: 3h
Pre-requisite: Any dense field
Use pointing mask to also measure mechanical center of masks

> Rotator center
  • Configure for imaging with N3.75, Ks
  • Preset to any dense star field at PA=0
  • Allow to collimate
  • Take image
  • Preset to same field at PA=180
  • Allow to collimate
  • Take image
  • Also do PA=90, 270
  • Repeat on at least two different fields

> Mean coords of matched stars gives rotator center > Mean coords of reference circle gives mechanical center

> Update the configuration file with the N1.80 rotator center coords.

Calibration of AGw transform Duration: 6h
See: http://wiki.lbto.org/twiki/bin/view/Commissioning/TransformationData

Field Aberration Measurement
See: http://wiki.lbto.org/twiki/bin/view/Commissioning/FieldAberrationMeasurementDetails

Verification of AGw transform Duration: 1h
Once transform is in place, take another set of data to verify.

Note: a repeated dice-5 pattern should be well-aligned with rows/columns

PM: Total system throughput and photometric ZPs: Duration: 7h
In all BB-filters
Pre-requisite: photometric conditions

### Standard with a range of colors observed at a range of airmass to
### derive ZP@airmass=1 and extinction, perhaps also color terms.

> Photometric zero points
  • Preset to series of photometric standards (AIRM 1.0…2.0)
  • Take imagepair through each BB filter (K, Ks, H, J, z)
  • Do only filters for which there is a good (and recent!) flatfield available.
  • Minimum ~5 samples per filter spread in airmass, 10 is better

PM: Field illumination: Duration: 3h
N3.75 camera due to new field lens
Pre-requisite: photometric conditions
3h should be enough to do a couple of filters
See /home/observer/LBTO/SciencePrep/Illumination for generic scripts

> Field Illumination
  • Configure for imaging in N3.75 camera
  • Select filter, target (any Persson standard, other fields?)
  • Test DIT, configure NDIT to allow collimation
  • Center source on detector
  • Run at least two scripts per filter:
    • Ddice5_100as_illum.scr (gets the extreme corners, +/-100")
    • D25x_illum1.scr (covers field, monitors extinction)
  • Repeat for all filters (eventually) BB and NB.

Ghosts, scattered light: Done? Duration: 2h
Not a must but nice add-on.
Pre-requisite: bright star
Bad weather filler

> Ghost check
  • Configure for imaging, J band, N3.75 camera
  • Point to bright star (any in M5 catalog, pref. not K giants)
  • Offset to put the source in the center of the top-left quadrant
  • Take a 7x4.0s exposure
  • Offset DETXY dy=-120" to place in bottom-left quadrant
  • Take a 7x4.0s exposure
  • Difference to check for ghosting, scattered light

PM: Spectrophotometry tests of the new detector and readout modes Duration: 3h
In principle not necessary since RON is known, but since spectroscopy is the dominant observing mode it might be wise to check.
Pre-requisite: spectrophot. std

> Specphot tests of new detector
  • Configure for specphot (wide slit)
  • Preset to any spectrophotometric standard
  • Take data over all major modes (gratings and bandpasses)
  • zJspec at high priority as it is new

Specphot test table
Camera Filter Grating Mask Notes
N1.80 clear K 210_zJHK 3_slit #3 10x30" wide slits

Run a complete science program

LUCI-to-AGw Focus Measurement Duration: 2h
Pre-requisite: STABLE temps, telescope in good thermal equilibrium, good (<0.7") seeing

Walter, can you outline how you did this?

Note: Can monitor the seeing from the guide cam during measurements to check for any systematic variations. The guide cam will always be in focus once collimation has converged.

> LUCI-to_AGw Focus Measurement

Still to add to the plan?

-- DavidThompson - 20 Aug 2013
Topic revision: r3 - 10 Sep 2013, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback